

********************************************************************
UARS WEEKLY STATUS REPORT - 19 JUNE 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE proceeds with its customary operating status, continuing to
accumulate celestial statistics during all obtainable solar observation
intervals and astral calibration analyses, whenever feasible.
Happy SOLSTICE to all!
***********
SUSIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
The MLS instrument has been powered off since June 1. MLS is scheduled to
begin north viewing operations on June 28.
***********
HRDI
***********
HRDI is currently in low power mode. HRDI is ready to begin day-side only
operations when the required ground software is available.
***********
WINDII
***********
WINDII WEEKLY REPORT FOR THE WEEK ENDING JUNE 19, 1998
WINDII is scheduled to be activated on Monday, June 22, and will be
concentrating on observations of polar mesospheric clouds.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no other problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
An orbit-adjust maneuver for the UARS spacecraft is being planned
for 12 August 1998. In preparation for the thruster firing on
12 August, small thruster "burns" for calibration are being
planned for 28 July 1998.
***********
CDHF
***********
The CDHF received a total of 395 CD-ROM requests as of 18-June-1998.
There were two orders received and shipped this week.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
27 days PEM (3.32) PROD (1379R-1366C=13L)
50 days HAL (4.80) PROD (L1) (1868R-442C-2F=1424L)
183 days HAL (4.80) PROD (L3) (1868R-288C=1580L)
603 days SOL (4.61) PROD (L1) (2646R-2066C=580L)
183 days SOL (4.61) PROD (L2-L3) (2349R-428C=1921L)
6 days CLA (7.00) ENG (6R-2C=4L)
Preventive Maintenance for the UARS CDHF systems is due in June. It is
scheduled and confirmed for Tuesday 06/23/98 in conjunction with the UARSDB
Full Table Maintenance.
UARS PCG will load April 1998 current day DAAC data in June 1998.
CDHF pending software upgrades: IDL 5.1, OpenIngres
***********
FOT
***********
REPORTING PERIOD
This report covers 06 June 1998 (Orbit 36798, GMT Day 157) through 12 June
1998 (Orbit 36902, GMT Day 163).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HRDI 161/2257 36872 Off (low power)
HALOE 163/2307 36902 Off
The beta angle is now decreasing toward a minimum angle of 0.0 degrees on
15 June 1998 (DOY 166).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 24.2 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.2 degrees C. Additional battery performance data for
each battery on 06 June 1998 and 12 June 1998 are provided in Appendix D.
The clock error ranged from -10.1 to +10.0 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = none
% OUT-OF-LIMITS = 06
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 03
% HRDI 01
% HALOE 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 02
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/146 - 2287.496748
Transponder B: 98/147 - 2287.499000
GROUND SYSTEM OPERATION
Two (2) operational support problem(s) occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/162 show the observatory data loss to be
20 hours, 43 minutes, 54.516 seconds (an increase of 17.888 seconds since
last report period). This is a 0.0352 percent data loss which equals a
99.9648 percent data capture for the mission.
The increase of 17.888 seconds occurred over a four day period:
7 4.480 seconds on Day 156 due to hit in data,
7 4.352 seconds on Day 159 due to hit in data,
7 5.440 seconds on Day 161 due to hit in data, and
7 3.616 seconds on Day 162 due to hit in data.
The latest release of NRT/quicklook processing software on UMPGAL for use
with UGPL was tested. For the first time, the software ran to completion
without error. However, the input files still have some peculiarities
(spurious data points at the end of the data set) and have not been
verified to contain valid data quality files.
FOT received a request from SUSIM Operations to perform a series of Gridded
Offset maneuvers on GMT Day 171 through 176.
FOT received requests from SOLSTICE operations to perform two Rmanually
generatedS stellar offset experiments on GMT Days 167 and 169.
FOT continued testing Database version 67 this week. Most of the PMON
logic has been tested, but other areas remain untested, specifically the
reduced ATC capabilities. However, the FOT cannot proceed until an AP disk
pack that containing corrected OBC dump comparison software is delivered.
Testing the command side of the I/P Net remains on hold. Testing will
begin when MSOCC provides a command loop interface from the POCC through
the I/P Net to the UARS Simulator.
OTHER SIGNIFICANT EVENTS
Continued testing of the upgrade to the OBC Power Monitor (PMON) flight
software. Ground System testing continued this week. Completion of
supporting documentation for PMON software uplink (system table
definitions, etc.) is in process.
Multiple simulations on the new AP software were conducted. However, in
the course of testing, a problem with the OBC Bank Dump capability was
observed. The test pack anomaly is associated with the default parameters
for the OBC bank dump directive. Previous versions of AP software worked
with the default parameters, but in this version, the parameters must be
explicitly declared. If all parameters are not specified, the ground
system will hang up. The problems was discussed with the programmer and
hopefully corrected software will be available next week.
Work continues on the UARS orbit adjust planning re-scheduled for 12 August
1998. Additional options for an Orbit Adjust are being planned in the
event the B/D thrusters can not be safely used. A status meeting was held
09 June 1998 to review progress of the task. Latch valve drive circuit
data from Hughes and Fairchild drawings were assembled for delivery to the
Propulsion Branch on 08 June 1998.
FOT continues to support CNMOS in their efforts to make necessary software
changes (MSOCC, CMS, DCF, etc.) to support the year 2000 rollover.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
The AP software programmer and FOT members met on 08 June 1998 to discussed
the problems that that occurred during last week with the AP software. AP
software programmer noted the problems and took an action to work the
problems.
A follow-up meeting to discuss the planned orbit adjust was held on 09 June
1998 with the Project Manager plus FDF and FOT representatives . The
Project Manager stated that the B/D attitude thrusters will probably not be
used, but may be refilled after Code 570 completes analysis of the proposed
FOT line refill procedure. The date for completion of the Code 570 study
is currently undefined.
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 15 June 1998.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
is targeted for 12 August 1998.
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 12 June 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
36798 157/0126 98-191 OBC Clk Adj fr 0.00 to 20.6 msec/day
36824 158/1910 98-192 OBC Clk Adj fr 20.6 to 0.00 msec/day
36872 161/2257 98-193 HRDI Power OFF
36892 163/0700 98-194 OBC Clk Adj fr 0.00 to 20.6 msec/day
36902 163/2307 98-195 HALOE Power OFF
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 12 June 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
WIEUTEMP Y-LO WINDII EU Temperature 36798-36902
HRPTSPBIMON Y-HI HRDI PST Pulse Bus Current Mon. 36805
HRPSEQBIMON Y-HI PSE Quiet Bus Current Monitor 36849,65-6,80,93
HRRADTMP Y-HI HRDI Radiometer Temp. 36874
HRIFESURVTMP Y-HI HRDI Inferometer Survival Temp 36874,78
QTSISAIF Y-HI SSPP Baseplate near SISA I/F 36884,89,93,95
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 12 June 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
36798 157/01:32 (06 Jun 98) from 00.0 to 20.6 msec/day
36824 158/19:24 (07 Jun 98) from 20.6 to 00.0 msec/day
36892 163/07:11 (12 Jun 98) from 00.0 to 20.6 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 12 June 1998
06 Jun 1998 - GMT Day 157 Beta = 38.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.2 0.000 0.0
2 +22.4/ -22.4 +6.11/+4.51 +29.6/-16.0 25.0 1.021 22.3
3 +78.4/ -11.2 +5.15/+3.23 +28.4/-16.4 25.0 1.025 21.6
12 Jun 1998 - GMT Day 163 Beta = 14.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.2 0.000 0.0
2 +22.4/ -11.2 +6.44/+5.15 +31.2/-15.2 24.2 1.022 23.1
3 +84.0/ -11.2 +5.15/+3.87 +30.0/-16.0 24.2 1.020 22.6
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
APPENDIX E
UARS ANOMALY REPORTS
Week ending 12 June 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-028 FOT 36848 160/0830 14170 Generic Late Acquisition
98-024 FOT 36887 163/0002 19991 Late Acq.-No XMTR On
Command in ATC
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-021 FOT MAY 1998 14170 GENERIC LATE ACQ
98=026 FOT 36680 149/03:12 TAC 4 CRASH
98-027 FOT 36698 150/08:19 TAC 4 CRASH
98-029 SUSIM 36765 154/20:36 BAD MICRO LOAD
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-028 Generic Late Acq for June 1998 (Open until 30 June 1998).
98-024 FOT 36887 163/0002 19991 Late Acq.-No XMTR On
Command in ATC
SEVNTFW24